1. a stone is dropped from the top of a tower. when it crosses a point 5 m below the top, another stone is let fall from a point 25m below the top. both stones reach the bottom of the tower simultaneously. find the height of the tower. g=10

Grade:11

1 Answers

Prudhvi teja
13 years ago

dear jadhayla

time taken for 25 ms=(2*25/9.8)0.5

time taken for 5 mis = (2*5/9.8)0.5

total time taken  = (2*h/9.8)0.5 = (2*25/9.8)0.5 + (2*5/9.8)0.5

total height = 52.306 m

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian